The Saunders Investment Bank has the following financing outstanding.

Debt:

46,000 bonds with a coupon rate of 5.1 percent and a current price quote of 105.3; the bonds have 15 years to maturity and a par value of $1,000. 16,100 zero coupon bonds with a price quote of 24.8, 30 years until maturity, and a par value of $10,000. Both bonds have semiannual compounding.

Preferred stock:

141,000 shares of 3.6 percent preferred stock with a current price of $87 and a par value of $100.

Common stock:

2,020,000 shares of common stock; the current price is $78 and the beta of the stock is 1.10.

Market:

The corporate tax rate is 21 percent, the market risk premium is 7.1 percent, and the risk-free rate is 3.1 percent.

What is the WACC for the company? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=78*2020000
=157560000
MV of Bond1=Par value*bonds outstanding*%age of par
MV of Bond1=1000*46000*1.053
=48438000
MV of Bond2=Par value*bonds outstanding*%age of par
MV of Bond2=10000*16100*0.248
=39928000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=87*141000
=12267000
MV of firm = MV of Equity + MV of Bond1+ MV of Bond 2+ MV of Preferred equity
=157560000+48438000+39928000+12267000
=258193000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 157560000/258193000
W(E)=0.6102
Weight of debt = MV of Bond/MV of firm
Weight of debt = 88366000/258193000
W(D)=0.3422
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 12267000/258193000
W(PE)=0.0475
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 3.1 + 1.1 * (7.1)
Cost of equity% = 10.91
Cost of debt
Bond1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =15x2
1053 =∑ [(5.1*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^15x2
                   k=1
YTM1 = 4.606740987
Bond2
                  K = Nx2
Bond Price   =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                   K =30x2
2480 =∑ [(0*10000/200)/(1 + YTM/200)^k]     +   10000/(1 + YTM/200)^30x2
                    k=1
YTM2 = 4.7
Firm cost of debt=YTM1*(MV bond1)/(MV bond1+MV bond2)+YTM2*(MV bond2)/(MV bond1+MV bond2)
Firm cost of debt=4.606740987*(48438000)/(48438000+39928000)+4.7*(48438000)/(48438000+39928000)
Firm cost of debt=4.65%
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 4.65*(1-0.21)
= 3.6735
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 3.6/(87)*100
=4.14
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=3.67*0.3422+10.91*0.6102+4.14*0.0475
WACC =8.11%